Calculate the expected solar radiative flux at our latitude (40°N) at solar noon (time of peak radiation) at three times of year: equinox, and both summer and winter solstice. (Please ignore the role of the atmosphere here – in other words calculate your answer as if there were no atmosphere altering the incoming radiation.)
Answer #1

H = 150 *(time - 12), where H is hour angle

Z = cos-1 (sinXsinY + cosXcosYcosH), where Z is the zenith angle, X is latitude and Y is the solar declination angle

Y at equinox is 00 , Y during summer solstice is 23.450 and Y during winter solstice is -23.450.

Solar radiative flux = ScosZ, where S is the solar constant 1000 W/m2

H = 150*(12-12) = 0, X = 400

During  equinox

Z = cos-1 (sin400sin00 + cos400cos00cos00) = cos-1 (0+cos400*1*1) = 400

Solar radiative flux = 1000*cos400 = 766.044443 W/m2

During summer solistice

Z = cos-1 (sin400sin23.450 + cos400cos23.450cos00) = cos-1 (0.25579645+0.70277507) = cos-1 (0.95857152) = 16.54999990

Solar radiative flux = 1000*cos16.54999990 = 958.57152 W/m2

During Winter solistice

Z = cos-1 (sin400sin(-23.450) + cos400cos(-23.450)cos00) = cos-1 (-0.255796449+0.70277507) = cos-1 (0.446978621) = 63.450

Solar radiative flux = 1000*cos63.450 = 446.978621 W/m2